# On a family of arithmetic series related to the Möbius function

### v5

@@ 1 Introduction and statements
${\mathcal{P}}:={\mathbb{P}}\cap\cup_{j\geqslant 1}\big]\sqrt{x_{j}},x_{j}]$
@@
$\liminf_{x\to\infty}\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{P}}\end{subarray}}{\mu(n)\omega(n)\over n}=-\log 2.$
This turns out to be a straightforward consequence of (2.4) infra.
@@
Uniformly for $2\leqslant y\leqslant\sqrt{x}$, we have
@@
It remains to estimate the contribution from ${\mathcal{Q}}_{y}:={\mathcal{P}}\smallsetminus{\mathcal{P}}_{y}$ to the sum (1.5), viz.
@@
Let us insert this into (2.10) and then into (2.11) keeping in mind the hypothesis $\log y\geqslant(\log_{2}x)^{b}$. Using the estimate ${\rm e}^{-J(s)}\asymp\min(|s|,1)$ $(\Re s\geqslant-1)$ proved in [HT93, lemma 2] in the form

### v6

@@ 1 Introduction and statements
${\mathcal{P}}:={\mathbb{P}}\cap\cup_{j\geqslant 1}\big]\sqrt{x_{j}},x_{j}]$ (1.6)
@@
$\liminf_{x\to\infty}\sum_{\begin{subarray}{c}n\leqslant x\\ P^{-}(n)\in{\mathcal{P}}\end{subarray}}{\mu(n)\omega(n)\over n}\leqslant-\log 2.$ (1.7)
This turns out to be a straightforward consequence of (2.4) infra: see Remark 2.3.
@@
Without any hypothesis on $\mathcal{P}$} and uniformly for $2\leqslant y\leqslant\sqrt{x}$, we have
@@
Remark 2.3.
To prove (1.7) observe that if $\mathcal{P}$} is defined by (1.6), then
$A^{-}\big(x_{j},\sqrt{x_{j}}\big)=A^{-}\big(x_{j},x_{j-1}\big)\ll\frac{\log x_{j-1}}{\log x_{j}}$
by Lemma 2.4, and
$\sum_{\begin{subarray}{c}n\leqslant x_{j}\\ P^{-}(n)\in{\mathcal{P}}\\ P^{-}(n)>\sqrt{x_{j}}\end{subarray}}\frac{\mu(n)\omega(n)}{n}=\sum_{\sqrt{x_{j}}<p\leqslant x_{j}}{-1\over p}=-\log 2+o(1)\qquad(j\to\infty).$
This implies (1.7) on selecting $\{x_{j}\}_{j=1}^{\infty}$ tending to infinity sufficiently fast.
To complete the proof of Theorem 1.1, it remains to estimate the contribution from ${\mathcal{Q}}_{y}:={\mathcal{P}}\smallsetminus{\mathcal{P}}_{y}$ to the sum (1.5), viz.
@@
(This is proved using the Korobov-Vinogradov zeo free region and accounts for the exponent $3/5$ in the definition of $L_{\varepsilon}(y)$.)
Let us insert (2.12) into (2.10) and then into (2.11) keeping in mind the hypothesis $\log y\geqslant(\log_{2}x)^{b}$. Using the estimate ${\rm e}^{-J(s)}\asymp\min(|s|,1)$ $(\Re s\geqslant-1)$ proved in [HT93, lemma 2] in the form